A Bibliography of Publications of Jorge Ricardo Paloschi

Jorge Ricardo Paloschi Imperial College Centre for Process Systems Engineering Prince Consort Rd. LONDON - SW7 2BY United Kingdom Tel: +44 171 594 6639 FAX: +44 171 594 6606 E-mail: j.paloschi@ic.ac.uk (Internet) 25 November 2011 Version 1.02

Abstract

This bibliography records publications of Jorge Ricardo Paloschi.

Title word cross-reference

algebraic [Pal94a]. algorithm [Pal94a]. bounded [Pal94b]. continuation [Pal94a]. equations [PP88a]. experiments [Pal94b]. factors [PP86]. homotopies [Pal94b]. hybrid [Pal94a]. implementation [PP88a]. invariant [PP88b].

LU [PP86].

methods [PP86].

Newton [PP86]. nonlinear [PP88a].

Numerical [Pal94b].

Quasi [PP86].

Scale [PP88b]. sets [PP88a]. solution [PP88b]. solve [Pal94a]. solving [PP88a]. systems [Pal95]. updating [PP86]. using [Pal94b].

References

Paloschi:1994:HCA

[Pal94a] J. R. Paloschi. A hybrid continuation algorithm to solve algebraic nonlinear equations. Comput. Chem. Engng., 18 Supl:S201–S209, 1994.

Paloschi:1994:N

[Pal94b] J. R. Paloschi. Numerical experiments using bounded homotopies to solve algebraic nonlinear equations. Proceedings of Escape 4 - Dublin - March 1994, 1994.

Paloschi:1995:BHS

[Pal95] J. R. Paloschi. Bounded homotopies to solve systems of algebraic nonlinear equations. Comput. Chem. Engng., To appear, 1995.

Paloschi:1986:ULF

[PP86] J. R. Paloschi and J. D. Perkins. The updating of LU factors in Quasi Newton methods. Comput. Chem. Engng., 10:241–247, 1986.

Paloschi:1988:IQN

[PP88a] J. R. Paloschi and J. D. Perkins.

An implementation of Quasi Newton methods for solving sets of nonlinear equations. Comput. Chem. Engng., 12:767–776, 1988.

Paloschi:1988:SIQ

[PP88b] J. R. Paloschi and J. D. Perkins.

Scale invariant Quasi Newton methods for the solution of nonlinear equations. Comput. Chem. Engng., 12:91–97, 1988.